ZoeyBiocth16:

(4x^2y^4)^3/2 can be simplified to which of the following? A) 8x^3y^6 B) 4x^5/2y C) 4xy D)32x^7/2y^11/2

ZoeyBiocth16:

i hope yall know what (^) means lol i have the answer i wanna make sure my work is right

Pixel:

what's your answer then

ZoeyBiocth16:

one way i did it was \[(4x ^{2}y ^{4})^{3/2} = 4 ^{3/2} (x ^{2}) ^{3/2} (y ^{4})^{3/2} simplified 4^{3/2 } = (\sqrt{4 )^{3 }}\]

ZoeyBiocth16:

then 8x^2 3/2 y^4 3/2 = 8x^3 y^6

ZoeyBiocth16:

thats what i got

Pixel:

i used a calculator

Pixel:

that took 1 minute and got A

ZoeyBiocth16:

lmao i didnt.... i didnt have one at the time

ZoeyBiocth16:

its still right tho lmao